OpenStudy (anonymous):

When using comparison theorem to determine whether an integral is convergent or divergent how do you know which function to use. For example, one specific questions asks x/sqrt(1+x^6) between 1 and infiniti (lets call this g(x)). Now in order to use comparison theory we have to find a function, f(x) which is greater or equal to g(x). Can I let f(x) =x which will always be greater than g(x)? Or is there some sort of trick to figure out which function to use?

OpenStudy (anonymous):

There is no special trick. Just like the comparison tests for infinity series. If you have a series: \[\sum_{k=1}^{\infty}\frac{1}{n}\] You can compare it to: \[\sum_{k=1}^{\infty}\frac{2}{n}\] because its just a tiny bit bigger. (It may not be practical) So x is always bigger. You know that integral converges. That means a smaller integral must also converge.

OpenStudy (anonymous):

should read "infinite series." in the first line.

OpenStudy (anonymous):

\[\lim_{t \rightarrow \inf}\int\limits_{1}^{t}xdx = (x^2)/2\ =\lim_{t \rightarrow \inf} (t^2)/2) -1/2\] This would be divergent? which would mean that g(x) would also be divergent correct?

OpenStudy (anonymous):

If a larger series diverges, that says nothing about a smaller series :P You would have to find a smaller integral that diverges to show that a larger integral diverges.

OpenStudy (anonymous):

Try this and go down to comparison tests. http://science.kennesaw.edu/~plaval/math2202/intimp.pdf

OpenStudy (anonymous):

I deleted the first one because it wasn't the right link.

OpenStudy (anonymous):

Well, I am confused the theorem in my book says if \[f(x) \ge g(x)\] and f(x) either diverges or converges then g(x) does the same.. I dont understand why I cant let the smaller one be g(x) and come up with an f(x) that is greater...

OpenStudy (zarkon):

that is not true

OpenStudy (anonymous):

Zarkon are you talking about his last post or mine?

OpenStudy (zarkon):

Misterchyme post

OpenStudy (zarkon):

this f(x)≥g(x) and f(x) either diverges or converges then g(x) does the same is not ture

OpenStudy (zarkon):

it is possible for f(x) to diverge but g(x) to converge

OpenStudy (anonymous):

I thought that was for series not improper integrals

OpenStudy (zarkon):

\[g(x)=\frac{1}{x^2}\] \[f(x)=x^2\] clearly \[g(x)\leq f(x)\] \[\int\limits_{1}^{\infty}g(x)dx\leq\infty\] but \[\int\limits_{1}^{\infty}f(x)dx=\infty\]

OpenStudy (anonymous):

because my book.. in comparison test for improper integrals says.. Suppose that f and g are continious functions with F(x)>g(x)>0 for x>a If f(x) converges then g(x) converges.. If g(x) diverges then f(x) diverges

OpenStudy (zarkon):

that is not what you have above

OpenStudy (anonymous):

Exactly. Notice the switch of the order on f(x) and g(x) from converge to diverge.

OpenStudy (anonymous):

Its the same concept I was getting at anyways. If f(x) is bigger and converges that FORCES the smaller integral to converge. If g(x) diverges and is a SMALLER integral then obviously the bigger one diverges as well.

OpenStudy (anonymous):

Remember, an integral is just an infinite series :P

OpenStudy (anonymous):

For some reason I am completely lost

OpenStudy (anonymous):

I am prob missing something really simple that is causing this confusion

OpenStudy (zarkon):

you know this correct? if \[g(x)\leq f(x)\] for all \[x\geq a\] then \[\int\limits_{a}^{b}g(x)dx\leq \int\limits_{a}^{b}f(x)dx\] or to extend it a little \[\int\limits_{a}^{\infty}g(x)dx\leq \int\limits_{a}^{\infty}f(x)dx\]

OpenStudy (zarkon):

the integral preserves the order

OpenStudy (anonymous):

Yea i get that, but I dont understand how/what function to use for comparison test.. Something just isnt clicking right now. How do I know wether to chose a function that is less than the original function or greater then is my dilemma

OpenStudy (zarkon):

a lot of the time you end up comparing the original function to \[\frac{1}{x^p}\]

OpenStudy (zarkon):

\[\int\limits_{a}^{\infty}\frac{1}{x^p}dx\] converges for p>1 and diverges otherwise

OpenStudy (zarkon):

I should specify that a>0 ;)

OpenStudy (anonymous):

well for instance my original function was \[\int\limits_{1}^{\infty} x/(\sqrt(1+x^6)\] How do I know whether to use f(x)=x which is \[\ge x/(\sqrt(1+x^6)\] or to use \[1/\sqrt(1+x^6) \le \ x/(\sqrt(1+x^6)\]

OpenStudy (zarkon):

\[\frac{x}{\sqrt{1+x^6}}\leq\frac{x}{\sqrt{x^6}}=\frac{x}{x^3}=\frac{1}{x^2}\]

OpenStudy (anonymous):

see, that stuff just doesnt pop out to me.. I see it now that you have pointed it out Lol.

OpenStudy (zarkon):

do you know now if it will converge or diverge based on what i did? the integral from 1 to infinity

OpenStudy (anonymous):

yes, it will converge because 2>1 correct? ( I hope so!)

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Well now that I know what to look for, I should be ok ( hopefully) with a little more practice. Thanks for your help!